OpenStudy (anonymous):

how would you write a slope-intercept equation with a using the points (8,-3) and (5,-2)?

OpenStudy (anonymous):

first find the slope between the two lines: slope = rise/run sound familiar? yes no? maybe so?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

good so can you solve for the slope for me? ya no?

OpenStudy (anonymous):

-1/3

OpenStudy (anonymous):

yep, good work

OpenStudy (anonymous):

now we know m=-1/3 lets get into making this an equation: y=mx+b is the standard form for 'slope intercept' plug in m y=(-1/3)x+b we now have to find the y-intercept, or 'b' so plug in one of the two points they give you into our equation y=(-1/3)x+b and solve for 'b'

OpenStudy (anonymous):

what do you mean plug in one of the two points?

OpenStudy (anonymous):

ok so I mean something like this: take the point (5,-2) and plug it into the equation we've developed: y=(-1/3)x+b becomes -2 = (-1/3)(5) + b then we just need to solve the equation for 'b'

OpenStudy (anonymous):

i'm not sure how to solve this equation

OpenStudy (anonymous):

thats ok, I can teach you, when ever we want to 'isolate' a term or 'solve for' a term, we need to get that term by it's self on it's own side of the equal sign. we are solving for 'b', so we want to make the equation so 'b' is the thing on the side the right side of the equal sign.

OpenStudy (anonymous):

so subtract b from each side then subtract -2 from each side

OpenStudy (anonymous):

you got the right idea, but i'll give you some more direction/ example

OpenStudy (anonymous):